how do you solve 2x+25=12
Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(48561)

2x+25=12
1st you subtract 25 from each side to get:
2x+25-25 = 12-25
2x=-13
2nd you divide each side by 2 to get:
x=-13/2
Cheers,
Stan H.

Answer by tutorcecilia(2144)

2x+25=12
2x+25-25=12-25
2x=-13
2x/2=-13/2
x = -13/2
.
Check by plugging (x=-13/2) back in the original equation:
2x+25=12
2(-13/2)+25=12 [Simplify]
-26/3+25=12
12=12 [Checks out]